OpenStudy (anonymous):

An electron and a proton are each placed at rest in a uniform electric field of magnitude 554 N/C. Calculate the speed of each particle 51.6 ns after being released.

OpenStudy (anonymous):

\[F=Eq=554*1.6*10^{-19}=ma=9.1*10^{-31}a\]

OpenStudy (anonymous):

\[s=\frac{1}{2}at^2=\frac{554*1.6*10^{-19}}{2*9.1*10^{-31}}(51.6*10^{-9})^2\]

OpenStudy (anonymous):

For a proton, replace 9.1*10^-31 with the proton's (kg) mass

OpenStudy (anonymous):

Do you get this?

OpenStudy (anonymous):

@riptidej ?

OpenStudy (anonymous):

i dont

OpenStudy (anonymous):

i'm sorry..i just never understood physics and i just dont get it

OpenStudy (anonymous):

Fair enough. OK- the Electric field strength at a point is just the force exerted on a unit charge. Is that OK?

OpenStudy (anonymous):

(btw I will NOT explain it as well as http://www.youtube.com/watch?v=OsWDUqJQcpk&feature=relmfu if you have the time)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i will take a look

OpenStudy (anonymous):

So E at a point is in the same direction as the force, but is the F/q: force divided by the charge on that object, as force in a way is useless in that you don't know how strong the electricity is unless you account for the strength of the charge it is moving

OpenStudy (anonymous):

So E=F/q F=E*q, which by Newton's 2nd law is also =ma

OpenStudy (anonymous):

Stop me if you have any question!

OpenStudy (anonymous):

ok i see that in my book..that formular..but when i used it..i got the wrong answer. so i must have messed up the formula

OpenStudy (anonymous):

So the acceleration (to do with velocity/ speed- what we are trying to find, so we're going to isolate it) is equal to =a=Eq/m

OpenStudy (anonymous):

Did you try plugging in the numbers above?

OpenStudy (anonymous):

Sorry, my above formula gives the displacement, not the velocity. It should be \[v=at\] So take away the 1/2 and the squared sign on the time

OpenStudy (anonymous):

So: Acceleration is how fast the thing speeds up per second, so its speed is acceleration*time v=at You know a=Eq/m You know t Substitute the actual values and work out V

OpenStudy (anonymous):

google the electron/proton's mass, charge etc.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

For the electron it turns out as 886400

OpenStudy (anonymous):

it says thats wrong

OpenStudy (anonymous):

I made some approximations of the mass/charge of the electron. It obviously wants it more precise.

OpenStudy (anonymous):

it says its off by 10%